OpenStudy (anonymous):

PLLLLLLLLLLLEAAAAAAAAASEEE help me find the 30th term of the following arithmetic sequence 35,29,23,17

OpenStudy (ilovecake):

175?

OpenStudy (anonymous):

can you please show work so i can understand

OpenStudy (ilovecake):

okay

OpenStudy (anonymous):

thank you

OpenStudy (ilovecake):

t1=1 d=6 tn=t1+(n-1)d tn=2 t1+(n-1)dt 30 =30-1)= 175

OpenStudy (ilovecake):

There..Hope that helped

OpenStudy (anonymous):

it did alot thank you

OpenStudy (ilovecake):

Your welcome